100 grams of water was produced from the complete combustion of naphthalene (C10H8).

Calculate for the initial volume of the fuel in liters. Assume that it is stored in a container

under a pressure of 1 atm and a temperature of 25°C.


1
Expert's answer
2021-01-28T05:37:22-0500

Solution:

Fuels such as petroleum and coal contain naphthalene.


The balanced chemical equation:

C10H8 + 12O2 → 10CO2 + 4H2O

According to the equatione above: n(C10H8) = n(H2O)/4


Moles of H2O = Mass of H2O / Molar mass of H2O

The molar mass of H2O is 18.015 g mol-1.

Hence,

n(H2O) = 100 g / 18.015 g mol-1 = 5.55 mol


n(C10H8) = n(H2O) / 4 = 5.55 mol / 4 = 1.3875 mol

n(C10H8) = 1.3875 mol


The Ideal Gas equation can be used:

PV = nRT


P = 1 atm

V = unknown

n = 1.3875 mol

R = 0.082 L atm K-1 mol-1

T = 25°C + 273.15 = 298.15 K


Thus,

V(fuel) = nRT/P

V(fuel) = (1.3875 mol × 0.082 L atm K-1 mol-1 × 298.15 K) / (1 atm) = 33.922 L = 33.9 L

V(fuel) = 33.9 L


Answer: The initial volume of the fuel is 33.9 liters.